OpenStudy (anonymous):

Find all real solutions of the equation. (If there is no real solution, enter NO REAL SOLUTION.) 25x^2 + 70x + 49 = 0 x = _____

OpenStudy (anonymous):

x=-7/5

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

is it positive and negative 7/5

OpenStudy (saifoo.khan):

only -7/5

jhonyy9 (jhonyy9):

-70 +/- radical (4900-4900) x_1,_2= ----------------------- = -7/5 50 x=-7/5